# EvoRL: A GPU-accelerated Framework for Evolutionary Reinforcement Learning

### v2

@@ 1. Introduction
Reinforcement Learning (RL) is a machine learning paradigm aimed at training agents for sequential decision-making tasks. Unlike supervised learning, which operates on static datasets, RL learns an agent’s policy through interactions with the environment, seeking to maximize cumulative rewards (returns). Advances in gradient-based optimization within deep learning have driven RL to achieve significant success across various domains, including board games (Silver et al., 2016, 2018), video games (Mnih et al., 2015, 2016; Vinyals et al., 2019), and robotic control (Lillicrap et al., 2016; Fujimoto et al., 2018; Haarnoja et al., 2018). Despite its successes, RL faces critical challenges due to the non-stationary nature of its training environment. These challenges include instability arising from sensitivity to hyperparameters and the need to carefully balance exploration and exploitation. Beyond RL, the Evolutionary Computation (EC) has emerged as a competitive alternative for policy search. EC leverages population-based methods, e.g., Evolutionary Algorithms (EAs), which utilize episodic returns as feedback signals and ensure diverse exploration through the population. However, EAs are often constrained by high sample complexity and the inefficiencies of their gradient-free, heuristic-driven optimization strategies, resulting in slow convergence rates.
@@ 2.1. Reinforcement Learning
Model-free reinforcement learning (RL) is generally framed under Markov Decision Process (MDP) assumption (Sutton and Barto, 2018), where the problem is defined by a tuple $(\mathcal{S},\mathcal{A},p,r)$. $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ is the transition probability function, and $r:\mathcal{S}\times\mathcal{A}\mapsto[r_{\text{min}},r_{\text{max}}]$ is the reward function after every transition. The goal is to find an optimal policy $\pi(a_{t}|s_{t})$ that maximizes the expected return: $\mathbb{E}_{\tau\sim p(\tau|\pi)}\sum_{t}^{|\tau|}r_{t}$, where $\tau$ is a trajectory generated by the policy $\pi$. In addition, a discount factor $\gamma\in[0,1]$ is introduced as an inductive bias [^2] Although, in the episodic tasks, the discount factor has been treated as a part of the environment previously, it is deemed as a hyperparameter of RL algorithms (Hessel et al., 2019; Amit et al., 2020; Grigsby et al., 2021) in recent research, where agents’ undiscounted returns are reported for evaluation..
@@ 2.3. Evolutionary Reinforcement Learning
Another prominent EvoRL paradigm employs EAs as meta-algorithms for RL, referred to as population-based AutoRL. A notable example is Population-Based Training (PBT) (Jaderberg et al., 2017), which proposes an online hyperparameter tuning method. PBT trains a population of models with varying hyperparameters, where each individual consists of both hyperparameters and network weights. After each training interval, PBT selects high-performing individuals to replace poorly performing ones, based on the meta objectives of each individual, such as the average evaluation return. Then it perturbes the hyperparameters of the replaced individuals. PBT can be viewed as a specific instance of EC (Petrenko et al., 2023; Shahid et al., 2024), incorporating mutation and evolutionary selection operators. For example, Bai and Cheng (2024) introduced a pairwise learning operator inspired by the Cooperative Swarm Optimizer (CSO) (Cheng and Jin, 2015) to replace the standard mutation operator in PBT. Similarly, Dushatskiy et al. (2023) extended PBT into multi-objective scenarios using the non-dominated ranking approach from NSGA-II (Deb et al., 2002). Furthermore, EAs have been used as black-box optimizers to approximate meta-gradients which are often intractable in RL scenarios (Tang and Choromanski, 2020), offering a promising direction for hyperparameter tuning in RL.
Beyond hyperparameter tuning, EC has also been employed to search for reward functions (Faust et al., 2019; Sapora et al., 2024) or loss functions (Houthooft et al., 2018; Lu et al., 2022) in RL algorithms. Other integration strategies include using EC as a mechanism for RL action selection (Ma et al., 2022). For a more comprehensive review of these approaches, we refer readers to recent surveys (Bai et al., 2023; Lin et al., 2024; Li et al., 2024a; Sigaud, 2023).
@@ 2.4. GPU-Accelerated Environments
The primary speed-up achieved by GPU-accelerated environments stems from their ability to run a massive number of environment instances in parallel. This contrasts with traditional RL algorithms, which typically utilize only a few parallel environments. As a result, special algorithmic designs are necessary to efficiently handle the large-scale input data generated by these environments and achieve fast convergence. While several algorithms, such as IMPALA (Espeholt et al., 2018) and Ape-X (Horgan et al., 2018), have been proposed to address this challenge, the problem remains an open area for future research. Nonetheless, we believe that integrating EC into RL presents a promising direction to tackle this challenge.
@@ 2.5. GPU-Accelerated Evolutionary Algorithms
GPU acceleration benefits several components of EAs, such as random noise generation and matrix operations, which achieve substantial speed-ups compared to traditional CPU-based implementations. Furthermore, GPU-based EAs can seamlessly integrate gradient-based operations, enabling hybrid approaches that combine evolutionary search with gradient-based optimization. Additionally, GPU-accelerated EAs benefit from problems implemented on GPUs, facilitating the development of end-to-end training pipelines that eliminate the overhead caused by frequent communication between CPUs and GPUs.
@@ 3.2. Computation Model
Besides hierarchical vectorization, we also apply JIT compilation throughout the entire training logic in Workflow.step(), which effectively optimizes the tiny or redundant operations in the computation graph of the training pipeline. By combining these two techniques and the modular design, EvoRL achieves highly reusable components that enable the development of efficient, scalable, and readable algorithms. This design not only maximizes the computational potential of modern hardware but also simplifies the implementation of complex training pipelines, making EvoRL an effective framework for diverse EvoRL research needs.
@@ 4. Platform
Figure 4. The average iteration time of OpenES across different implementations. RLlib utilizes a distributed pipeline on CPUs, whereas ours employs an end-to-end pipeline on a single GPU, achieving a maximum speed-up of over 60x.
(a) CEM-RL
(b) PBT
Figure 5. The average iteration time of (a) CEM-RL and (b) PBT in different implementations. For CEM-RL, the fastpbrl implementation encounters out-of-memory issues when the population size exceeds 256, whereas our implementation achieves a 5–9x speed-up over the official implementation and a peak speed-up of 1.45x to fastpbrl. For PBT, we employ PPO as the underlying on-policy RL algorithm and SAC as the underlying off-policy RL algorithm. Due to memory constraints, the population sizes for PBT in RLlib and fastpbrl are also limited. Under these conditions, our implementation achieves a maximum speed-up of 30–40x.
@@ 5. Experiments
In this section, we comprehensively evaluate the performance, scalability, and efficiency of various algorithms implemented in EvoRL. Our experiments focus on continuous robotic locomotion tasks in Brax (Freeman et al., 2021), a GPU-accelerated physics engine. Since these tasks are rigorous replacements for the original CPU-based counterpart on Gym (Brockman et al., 2016), they serve as a robust benchmark for assessing EvoRL’s ability to leverage GPU acceleration and handle large-scale computational workloads, ensuring fair comparisons with other state-of-the-art implementations.
@@ 5.1. Acceleration and Scalability Performance
EvoRL constructs an end-to-end training pipeline for various algorithms, optimized for execution on modern accelerators such as GPUs. To evaluate the efficiency of these pipelines, we measure the average training time across different configurations and compare the results to prior implementations. For consistency, we use the Swimmer environment from Brax (Freeman et al., 2021), which features fixed-length episodes (1000 timesteps). This choice eliminates potential variability due to agent performance across implementations, ensuring fair comparisons. All experiments are conducted on a single machine equipped with dual Intel Xeon Gold 6132 CPUs (56 logical cores), 128 GiB RAM, and an Nvidia RTX 3090 GPU with 24 GiB VRAM. Detailed training settings and hyperparameters are provided in Appendix B.
@@ 5.1.1. Performance of EAs for Policy Search
We evaluate the performance of EvoRL’s OpenES implementation against the RLlib implementation [^4] https://docs.ray.io/en/latest/rllib, which uses the CPU-based distributed framework Ray. In RLlib, a large shared noise table is pre-built at the start of training to efficiently generate noise for new populations (Salimans et al., 2017). In contrast, EvoRL’s default implementation generates noise on-the-fly during training. To ensure fair comparisons, we also test a variant of EvoRL that incorporates a pre-built noise table.
@@ 5.1.2. Performance of ERL Algorithms
As shown in Fig. 5(a), EvoRL’s fully GPU-accelerated pipeline eliminates this communication overhead by performing rollouts, RL updates, and environment simulations entirely on the GPU. This design achieves a 5–9x speed-up compared to the official implementation and 1.45x speed-up to fastpbrl before its out-of-memory issues. EvoRL demonstrates robust scalability with increased population sizes, further emphasizing its efficiency in large-scale ERL tasks.
@@ 5.1.3. Performance of Population-based AutoRL Algorithms
In contrast, EvoRL executes the entire PBT pipeline on the GPU, including rollouts, RL updates, and a GPU VRAM-based shared replay buffer. Results, shown in Fig. 5(b), highlight that EvoRL achieves a maximun speed-up of 30-40x compared to both Ray and fastpbrl implementations. Furthermore, EvoRL avoids memory limitations observed in competing frameworks, supporting larger populations and more computationally intensive configurations.
@@ 5.2. Use Cases
The versatility and modular design of EvoRL enable it to support a wide range of algorithms, making it a valuable platform for EvoRL research. In this section, we demonstrate the applicability and performance of EvoRL through three key use cases, showcasing EvoRL’s ability to handle diverse algorithmic paradigms on the GPU-accelerated pipelines.
@@ 5.2.1. Use Case of EAs for Policy Search
As presented in Fig. 6, the results reveal that larger networks tend to achieve faster convergence but do not necessarily lead to superior final performance. For example, performance degradation is observed with $[256,256]$ networks in the Ant environment. Similarly, in the HalfCheetah environment, $[256,256]$ networks yield lower returns than $[16,16]$ networks when using ARS and VanillaES. Increasing the number of episodes for fitness evaluation improves the accuracy of fitness estimates in noisy environments. However, this improvement might come at the cost of reduced exploration by fitness noise, which may result in suboptimal solutions in certain cases. These findings highlight the trade-offs between network size and evaluation noise in evolutionary policy search. In addition, Table. 1 provides the best average return across different EAs on these environments. ARS and OpenES are two advanced EAs, archiving better performance for policy search.
@@ 5.2.2. Use Case of ERL Algorithms
The original implementations dynamically scale the number of RL updates per iteration based on the total number of timesteps collected—either from the current iteration for CEM-RL or from all iterations for ERL. However, dynamically scaling updates with cumulative timesteps results in escalating computational costs in later iterations. To address this, we modify ERL to align the number of RL updates per iteration with the timesteps sampled from the current iteration, similar to CEM-RL. Additionally, we introduce a fixed number of RL updates per iteration, decoupling the updates from population size and agent performance (e.g., episode length). This adjustment significantly reduces computational overhead, particularly for larger population sizes, while maintaining scalability.
@@ 6. Related Work
Limited libraries have explored EvoRL-specific functionalities. Examples include RLlib (Liang et al., 2018) and fastpbrl (Flajolet et al., 2022), which support PBT for hyperparameter optimization. fastpbrl also implements a few ERL algorithms, such as CEM-RL (Pourchot and Sigaud, 2019) and DvD (Parker-Holder et al., 2020). Similarly, Lamarckian (Bai et al., 2022) provides a platform for EvoRL, including EAs and PBT algorithms. However, these libraries are designed for CPU-based environments and rely on hybrid pipeline architectures to orchestrate heterogeneous devices. For example, RLlib and fastpbrl use multiprocessing to distribute rollout processes across multiple CPU cores, requiring manual asynchronous scheduling and signal control to manage the training pipeline efficiently. These architectures increase complexity and introduce communication overhead between CPUs and GPUs.

### v3

@@ 1. Introduction
Reinforcement Learning (RL) is a machine learning paradigm aimed at training agents for sequential decision-making tasks. Unlike supervised learning, which operates on static datasets, RL learns an agent’s policy through interactions with the environment, seeking to maximize cumulative rewards (returns). Advances in gradient-based optimization within deep learning have driven RL to achieve significant success across various domains, including board games (Silver et al., 2016, 2018), video games (Mnih et al., 2015, 2016; Vinyals et al., 2019), and robotic control (Lillicrap et al., 2016; Fujimoto et al., 2018; Haarnoja et al., 2018). Despite its successes, RL faces critical challenges due to the non-stationary nature of its training data source. These challenges include instability arising from sensitivity to hyperparameters and the need to carefully balance exploration and exploitation. Beyond RL, the Evolutionary Computation (EC) has emerged as a competitive alternative for policy search. EC leverages population-based methods, e.g., Evolutionary Algorithms (EAs), which utilize episodic returns as feedback signals and ensure diverse exploration through the population. However, EAs are often constrained by high sample complexity and the inefficiencies of their gradient-free, heuristic-driven optimization strategies, resulting in slow convergence rates.
@@ 2.1. Reinforcement Learning
Model-free reinforcement learning (RL) is generally framed under the Markov Decision Process (MDP) assumption (Sutton and Barto, 2018), where the problem is defined by a tuple $(\mathcal{S},\mathcal{A},p,r)$. $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ is the transition probability function, and $r:\mathcal{S}\times\mathcal{A}\mapsto[r_{\text{min}},r_{\text{max}}]$ is the reward function after every transition. The goal is to find an optimal policy $\pi(a_{t}|s_{t})$ that maximizes the expected return: $\mathbb{E}_{\tau\sim p(\tau|\pi)}\sum_{t}^{|\tau|}r_{t}$, where $\tau$ is a trajectory generated by the policy $\pi$. In addition, a discount factor $\gamma\in[0,1]$ is introduced as an inductive bias [^2] Although, in the episodic tasks, the discount factor has been treated as a part of the environment previously, it is deemed as a hyperparameter of RL algorithms (Hessel et al., 2019; Amit et al., 2020; Grigsby et al., 2021) in recent research, where agents’ undiscounted returns are reported for evaluation..
@@ 2.3. Evolutionary Reinforcement Learning
Another prominent EvoRL paradigm employs EAs as meta-algorithms for RL, referred to as population-based AutoRL. A notable example is Population-Based Training (PBT) (Jaderberg et al., 2017), which proposes an online hyperparameter tuning method. PBT trains a population of models with varying hyperparameters, where each individual consists of both hyperparameters and network weights. After each training interval, PBT selects high-performing individuals to replace poorly performing ones based on the meta objectives of each individual, such as the average evaluation return. Then it perturbs the hyperparameters of the replaced individuals. PBT can be viewed as a specific instance of EC (Petrenko et al., 2023; Shahid et al., 2024), incorporating mutation and evolutionary selection operators. For example, Bai and Cheng (2024) introduced a pairwise learning operator inspired by the Cooperative Swarm Optimizer (CSO) (Cheng and Jin, 2015) to replace the standard mutation operator in PBT. Similarly, Dushatskiy et al. (2023) extended PBT into multi-objective scenarios using the non-dominated ranking approach from NSGA-II (Deb et al., 2002). Furthermore, EAs have been used as black-box optimizers to approximate meta-gradients, which are often intractable in RL scenarios (Tang and Choromanski, 2020), offering a promising direction for hyperparameter tuning in RL.
Beyond hyperparameter tuning, EC has also been employed to search for reward functions (Faust et al., 2019; Sapora et al., 2024; Onori et al., 2024) or loss functions (Houthooft et al., 2018; Lu et al., 2022) in RL algorithms. Other integration strategies include using EC as a mechanism for RL action selection (Ma et al., 2022). For a more comprehensive review of these approaches, we refer readers to recent surveys (Bai et al., 2023; Lin et al., 2024; Li et al., 2024a; Sigaud, 2023).
@@ 2.4. GPU-Accelerated Environments
A viable solution to eliminate CPU-GPU bottlenecks in RL training is to implement simulations directly on GPUs, which are naturally suited for high-throughput parallel simulations. Liang et al. (2018b) firstly demonstrated the promising speed-up achieved by GPU-accelerated robotic simulations in RL. Expanding on this concept, Isaac Gym (Makoviychuk et al., 2021) introduced an end-to-end GPU-based RL training pipeline, effectively eliminating the communication overhead between CPUs and GPUs. Compared to traditional CPU-based simulators, Isaac Gym achieves a 100x–1000x speed-up in training. Several subsequent studies (Petrenko et al., 2023; Handa et al., 2023; Shahid et al., 2024) have further validated the effectiveness of this end-to-end training pipeline using Isaac Gym. More recently, Genesis (Authors, 2024) introduced even more efficient and universal physics engines with GPU support, further advancing the development of GPU-accelerated RL environments. Besides, Mujoco playground (Zakka et al., 2025) has recently been released for GPU-accelerated environments for a comprehensive study in robot learning.
@@
The primary speed-up achieved by GPU-accelerated environments stems from their ability to run a massive number of environment instances in parallel. This contradicts traditional RL algorithms, which typically utilize only a few parallel environments. As a result, special algorithmic designs are necessary to efficiently handle the large-scale input data generated by these environments and achieve fast convergence. While several algorithms, such as IMPALA (Espeholt et al., 2018) and Ape-X (Horgan et al., 2018), have been proposed to address this challenge, the problem remains an open area for future research. Nonetheless, we believe that integrating EC into RL presents a promising direction to tackle this challenge.
@@ 2.5. GPU-Accelerated Evolutionary Algorithms
GPU acceleration benefits several components of EAs, such as random noise generation and matrix operations, which achieve substantial speed-ups compared to traditional CPU-based implementations. Furthermore, GPU-based EAs can seamlessly integrate gradient-based operations, enabling hybrid approaches that combine evolutionary search with gradient-based optimization (Gangwani and Peng, 2018). Additionally, GPU-accelerated EAs benefit from problems implemented on GPUs, facilitating the development of end-to-end training pipelines that eliminate the overhead caused by frequent communication between CPUs and GPUs.
@@ 3.2. Computation Model
The hierarchical vectorization architecture plays a critical role in supporting efficient parallelism in various algorithmic components from different types of algorithms, as illustrated in Fig. 3. For example, evolutionary algorithms (EAs) require simultaneous evaluation of all agents in the population to compute fitness scores, where each agent interacts with multiple parallel environment copies to collect episode returns. This necessitates a two-level vectorization scheme to manage both agent-level and environment-level parallelism. In the case of CEM-RL, an even more complex vectorization strategy is employed in multiple dimensions. A population of agents is initialized, and at each iteration, half agents are randomly selected and updated using off-policy reinforcement learning (RL). After this, all agents undergo an evaluation to perform an EA update, and their batched trajectories are stored in a replay buffer for subsequent off-policy RL updates. Population-Based Training (PBT) further extends the need for hierarchical vectorization, as it orchestrates multiple concurrent workflows. It also requires parallel evaluations to compute meta-objectives (e.g., the average episode return) for evolutionary tuning of hyperparameters.
Besides hierarchical vectorization, we also apply JIT compilation throughout the entire training logic in Workflow.step(), which effectively optimizes the tiny or redundant operations in the computation graph of the training pipeline. By combining these two techniques, along with the modular design, EvoRL achieves highly reusable components that enable the development of efficient, scalable, and readable algorithms. This design not only maximizes the computational potential of modern hardware but also simplifies the implementation of complex training pipelines, making EvoRL an effective framework for diverse EvoRL research needs.
@@ 5. Experiments
In this section, we comprehensively evaluate the performance, scalability, and efficiency of various algorithms implemented in EvoRL. Our experiments focus on continuous robotic locomotion tasks in Brax (Freeman et al., 2021), a GPU-accelerated physics engine. Since these tasks are good replacements for the original CPU-based counterpart on Gym (Brockman et al., 2016), they serve as a robust benchmark for assessing EvoRL’s ability to leverage GPU acceleration and handle large-scale computational workloads, ensuring fair comparisons with other state-of-the-art implementations.
@@ 5.1. Acceleration and Scalability Performance
EvoRL constructs an end-to-end training pipeline for various algorithms, optimized for execution on modern accelerators such as GPUs. To evaluate the efficiency of these pipelines, we measure the average training time across different configurations and compare the results to prior implementations. For consistency, we use the Swimmer environment from both Gym (Brockman et al., 2016) and Brax (Freeman et al., 2021), which features fixed-length episodes (1000 timesteps). This choice eliminates potential variability due to agent performance across implementations, ensuring fair pipeline-level comparisons. All experiments are conducted on a single machine equipped with dual Intel Xeon Gold 6132 CPUs (56 logical cores), 128 GiB RAM, and an Nvidia RTX 3090 GPU with 24 GiB VRAM. Detailed training settings and hyperparameters are provided in Appendix B. In addition, Table 1 summarizes the essential training pipeline differences across different libraries.
Table 1. Training pipeline implementations on different libraries.
Library Environment Agent Update Rollout
EvoRL Brax Swimmer GPU GPU Vectorization
RLlib Gym Swimmer-v4 CPU Multiprocess (CPU)
(a) EA Implementation
Library Environment Agent Update Rollout
EvoRL Brax Swimmer GPU Vectorization GPU Vectorization
Official Gym Swimmer-v4 For-loop (GPU) For-loop (Sinlge CPU)
fastpbrl Gym Swimmer-v4 GPU Vectorization Multiprocess (CPU)
(b) CEM-RL Implementation
Library Environment Agent Update Rollout
EvoRL Brax Swimmer GPU Vectorization GPU Vectorization
fastpbrl Gym Swimmer-v4 GPU Vectorization Multiprocess (CPU)
RLLib (PBT) Gym Swimmer-v4 GPU Multiprocess Multiprocess (CPU)
(c) PBT Implementation
@@ 5.1.1. Performance of EAs for Policy Search
Figure 4. The average iteration time of OpenES across different implementations. RLlib utilizes a distributed pipeline on CPUs, whereas ours employs an end-to-end pipeline on a single GPU, achieving a maximum speed-up of over 60x in the RLlib range.
(a) CEM-RL
(b) PBT
Figure 5. The average iteration time of (a) CEM-RL and (b) PBT in different implementations. For CEM-RL, the fastpbrl implementation encounters out-of-memory issues when the population size exceeds 256, whereas our implementation achieves a 5–9x speed-up over the official implementation and a peak speed-up of 1.45x over fastpbrl. For PBT, we employ PPO as the underlying on-policy RL algorithm and SAC as the underlying off-policy RL algorithm. Due to memory constraints, the population sizes for PBT in RLlib and fastpbrl are also limited. Under these conditions, our implementation achieves a maximum speed-up of 30–40x.
We evaluate the performance of EvoRL’s OpenES implementation against the RLlib implementation [^4] https://docs.ray.io/en/latest/rllib, which uses the CPU-based distributed framework Ray. Their implementation uses a full CPU-based architecture, including rollout and agent update. Besides, in RLlib, a large shared noise table is pre-built at the start of training to efficiently generate noise for new populations (Salimans et al., 2017). In contrast, EvoRL’s default implementation generates noise on-the-fly during training. To ensure fair comparisons, we also test a variant of EvoRL that incorporates a pre-built noise table.
@@ 5.1.2. Performance of ERL Algorithms
As shown in Fig. 5(a), EvoRL’s fully GPU-accelerated pipeline eliminates this communication overhead by performing rollouts, RL updates, and environment simulations entirely on the GPU. This design achieves a 5–9x speed-up compared to the official implementation and a 1.45x speed-up to fastpbrl before its out-of-memory issues. EvoRL demonstrates robust scalability with increased population sizes, further emphasizing its efficiency in large-scale ERL tasks.
@@ 5.1.3. Performance of Population-based AutoRL Algorithms
In contrast, EvoRL executes the entire PBT pipeline on the GPU, including rollouts, RL updates, and a GPU VRAM-based shared replay buffer. Results, shown in Fig. 5(b), highlight that EvoRL achieves a maximum speed-up of 30-40x compared to both Ray and fastpbrl implementations. Furthermore, EvoRL avoids memory limitations observed in competing frameworks, supporting larger populations and more computationally intensive configurations.
@@ 5.2. Use Cases
The versatility and modular design of EvoRL enable it to support a wide range of algorithms, making it a valuable platform for EvoRL research. In this section, we demonstrate the applicability of EvoRL through three key use cases, showcasing EvoRL’s ability to handle diverse algorithmic paradigms on the GPU-accelerated pipelines.
@@ 5.2.1. Use Case of EAs for Policy Search
As presented in Fig. 6, the results reveal that larger networks tend to achieve faster convergence but do not necessarily lead to superior final performance. For example, performance degradation is observed with $[256,256]$ networks in the Ant environment. Similarly, in the HalfCheetah environment, $[256,256]$ networks yield lower returns than $[16,16]$ networks when using ARS and VanillaES. Increasing the number of episodes for fitness evaluation improves the accuracy of fitness estimates in noisy environments. However, this improvement might come at the cost of reduced exploration by fitness noise, which may result in suboptimal solutions in certain cases. These findings highlight the trade-offs between network size and evaluation noise in evolutionary policy search. In addition, Table. 2 provides the best average return across different EAs on these environments. ARS and OpenES are two advanced EAs, achieving better performance for policy search.
@@ 5.2.2. Use Case of ERL Algorithms
The original implementations dynamically scale the number of RL updates per iteration based on the total number of timesteps collected – either from the current iteration for CEM-RL or from all iterations for ERL. However, dynamically scaling updates with cumulative timesteps results in escalating computational costs in later iterations. To address this, we modify ERL to align the number of RL updates per iteration with the timesteps sampled from the current iteration, similar to CEM-RL. Additionally, we introduce a fixed number of RL updates per iteration, decoupling the updates from population size and agent performance (e.g., episode length). This adjustment significantly reduces computational overhead, particularly for larger population sizes, while maintaining scalability.
@@ 6. Related Work
Limited libraries have explored EvoRL-specific functionalities. Examples include RLlib (Liang et al., 2018a) and fastpbrl (Flajolet et al., 2022), which support PBT for hyperparameter optimization. fastpbrl also implements a few ERL algorithms, such as CEM-RL (Pourchot and Sigaud, 2019) and DvD (Parker-Holder et al., 2020). Similarly, Lamarckian (Bai et al., 2022) provides a platform for EvoRL, including EAs and PBT algorithms; and Pearl (Tangri et al., 2022) provides EC and ERL supports. However, these libraries are designed for CPU-based environments and rely on hybrid pipeline architectures to orchestrate heterogeneous devices. For example, RLlib and fastpbrl use multiprocessing to distribute rollout processes across multiple CPU cores, requiring manual asynchronous scheduling and signal control to manage the training pipeline efficiently. These architectures increase complexity and introduce communication overhead between CPUs and GPUs.